two complementary angles differ by 18 degrees. find the angles
Answer by TutorDelphia(193)

Complementary angles add up to 90.
We have two angles the larger and the smaller.
l-18=s
and
l+s=90
Since l-18=s we can sub in l-18 for s
l+l-18=90
2l=108
l=54
sub 54 in for l
54-18=s
s=36